OpenStudy (anonymous):

Write the point-slope form of an equation for a line that passes through the point with the given slope. (6, 3), slope = –1

OpenStudy (anonymous):

The point-slope form of the equation of a line is \[y-y _{1}=m(x-x _{1})\] then, the solution is \[y-3=-1(x-6)\]

OpenStudy (anonymous):

y + 3 = –1(x – 6)

OpenStudy (anonymous):

why that answer? the point is (6,3) or (6,-3)??

OpenStudy (anonymous):

y – 3 = (x – 6)

OpenStudy (anonymous):

Do you have a set of options?

OpenStudy (anonymous):

a. y + 3 = –1(x – 6) c. y – 3 = –1(x + 6) b. y – 3 = (x – 6) d. y – 3 = –1(x – 6)

OpenStudy (anonymous):

the answer is D. Check the picture. :) and check my first answer